# MeGA: Merging Multiple Independently Trained Neural Networks Based on Genetic Algorithm

### v3

@@ 1 Introduction
In this paper, we introduce our genetic algorithm-based method, MeGA, for merging the weights of multiple trained CNN models. Experiments on the CIFAR-10 dataset [18] demonstrate improvements in test accuracy compared to individual models and traditional techniques. Rather than focusing on comparative analysis, this research emphasizes the methodology of effectively merging models to leverage their collective strengths. Our results highlight the potential of genetic algorithms in optimizing neural network weight fusion, providing a scalable solution for integrating multiple pre-trained models in various deep learning applications
@@ 4.2 CIFAR-10 Results
Datasets Model (# of Params.) Baseline. Baseline Test Acc. Final Test Acc. # of classes (Random Seed) (Single Model) (Merged Model) ResNet 56 (0.86M) [12] Model 1 (56) 0.809 0.822 Model 2 (57) 0.793 ResNet 110 (1.78M) [12] Model 1 (56) 0.801 0.816 Model 2 (57) 0.814 ResNet 152 (3.51M) [12] Model 1 (56) 0.783 0.819 CIFAR-10 [18] Model 2 (57) 0.781 (10) Xception (22.96M) [12] Model 1 (56) 0.716 0.754 Model 2 (57) 0.730 DenseNet 121 (7.04M) [12] Model 1 (56) 0.719 0.742 Model 2 (57) 0.728 DenseNet 169 (12.65M) [12] Model 1 (56) 0.735 0.753 Model 2 (57) 0.731
@@
4.3 Extended Experiments for Multi-Models
In this section, we illustrate the application of our genetic algorithm-based weight merging method to combine multiple neural network models as depicted in Figure 3.
@@ 4.3 Extended Experiments for Multi-Models
Dataset Model (# of Params) Baseline. Baseline Test Acc. Final Test Acc. # of classes (Random Seed) (Single Model) (Merged Model) ResNet 56 (0.86M) [12] Model 1 (44) 0.804 0.822 Model 2 (45) 0.747 Model 3 (46) 0.788 Model 4 (47) 0.774 Model 5 (48) 0.806 Model 6 (49) 0.786 Model 7 (50) 0.749 Model 8 (51) 0.787 ResNet 110 (1.78M) [12] Model 1 (44) 0.782 0.824 Model 2 (45) 0.814 Model 3 (46) 0.744 CIFAR-10 [18] Model 4 (47) 0.800 (10) Model 5 (48) 0.801 Model 6 (49) 0.787 Model 7 (50) 0.781 Model 8 (51) 0.758 ResNet 152 (3.51M) [12] Model 1 (44) 0.775 0.815 Model 2 (45) 0.783 Model 3 (46) 0.781 Model 4 (47) 0.747 Model 5 (48) 0.771 Model 6 (49) 0.767 Model 7 (50) 0.772 Model 8 (51) 0.797
The hierarchical merging process starts by training eight models on the CIFAR-10 dataset [12, 18]. These models are paired and merged using a genetic algorithm, resulting in four intermediate models. These intermediate models are further merged into two higher-level models, which are finally merged into a single, robust model. This process ensures the final model combines the strengths of all eight original models. The results in Table 2 demonstrate the effectiveness of our genetic algorithm-based weight merging method. For ResNet 56 models, baseline test accuracies ranged from 0.747 to 0.806, improving to 0.822 after merging. ResNet 110 models had baseline accuracies between 0.744 and 0.814, with the merged model achieving 0.824. For ResNet 152 models, baseline accuracies ranged from 0.747 to 0.783, and the merged model reached 0.815. Despite the larger complexity of ResNet 152, our method effectively enhanced model performance.

### v4

@@ 1 Introduction
In this paper, we introduce our genetic algorithm-based method, MeGA, for merging the weights of multiple trained CNN models. Experiments on the CIFAR-10 dataset [18] demonstrate improvements in test accuracy compared to individual models and traditional techniques. Rather than focusing on comparative analysis, this research emphasizes the methodology of effectively merging models to leverage their collective strengths. Our results highlight the potential of genetic algorithms in optimizing neural network weight fusion, providing a scalable solution for integrating multiple pre-trained models in various deep learning applications. Additionally, we demonstrate that it is possible to merge the weights of neural networks that have been initialized differently and trained independently, further underscoring the flexibility and robustness of our proposed method.
@@ 4.2 CIFAR-10 Results
Datasets Model (# of Params.) Baseline. Baseline Test Acc. Merging Method Final Test Acc. # of classes (Random Seed) (Single Model) (Merged Model) ResNet 56 (0.86M) [12] Model 1 (56) 0.809 Weight Average 0.010 Model 2 (57) 0.793 MeGA (Ours) 0.822 ResNet 110 (1.78M) [12] Model 1 (56) 0.801 Weight Average 0.010 Model 2 (57) 0.814 MeGA (Ours) 0.816 ResNet 152 (3.51M) [12] Model 1 (56) 0.783 Weight Average 0.010 CIFAR-10 [18] Model 2 (57) 0.781 MeGA (Ours) 0.819 (10) Xception (22.96M) [12] Model 1 (56) 0.716 Weight Average 0.010 Model 2 (57) 0.730 MeGA (Ours) 0.754 DenseNet 121 (7.04M) [12] Model 1 (56) 0.719 Weight Average 0.010 Model 2 (57) 0.728 MeGA (Ours) 0.742 DenseNet 169 (12.65M) [12] Model 1 (56) 0.735 Weight Average 0.010 Model 2 (57) 0.731 MeGA (Ours) 0.753
@@
The reason why the weight averaging method resulted in poorer performance is that the networks were initialized differently and trained independently. This led to discrepancies in the learned weights, making simple averaging ineffective. This highlights the importance of considering the initialization and training processes when merging models.
4.3 Extended Experiments for Multi-Models
In this section, we illustrate the application of our genetic algorithm-based weight merging method to combine multiple neural network models as depicted in Figure 3. The hierarchical merging process starts by training eight models on the CIFAR-10 dataset [12, 18]. These models are paired and merged using a genetic algorithm, resulting in four intermediate models. These intermediate models are further merged into two higher-level models, which are finally merged into a single, robust model.
@@ 4.3 Extended Experiments for Multi-Models
Dataset Model (# of Params) Baseline. Baseline Test Acc. Merging Method Final Test Acc. # of classes (Random Seed) (Single Model) (Merged Model) ResNet 56 (0.86M) [12] Model 1 (44) 0.804 Weight Average 0.010 Model 2 (45) 0.747 Model 3 (46) 0.788 Model 4 (47) 0.774 Model 5 (48) 0.806 MeGA (Ours) 0.822 Model 6 (49) 0.786 Model 7 (50) 0.749 Model 8 (51) 0.787 ResNet 110 (1.78M) [12] Model 1 (44) 0.782 Weight Average 0.010 Model 2 (45) 0.814 Model 3 (46) 0.744 CIFAR-10 [18] Model 4 (47) 0.800 (10) Model 5 (48) 0.801 MeGA (Ours) 0.824 Model 6 (49) 0.787 Model 7 (50) 0.781 Model 8 (51) 0.758 ResNet 152 (3.51M) [12] Model 1 (44) 0.775 Weight Average 0.010 Model 2 (45) 0.783 Model 3 (46) 0.781 Model 4 (47) 0.747 Model 5 (48) 0.771 MeGA (Ours) 0.815 Model 6 (49) 0.767 Model 7 (50) 0.772 Model 8 (51) 0.797
This process ensures the final model combines the strengths of all eight original models. The results in Table 2 demonstrate the effectiveness of our genetic algorithm-based weight merging method. For ResNet 56 models, baseline test accuracies ranged from 0.747 to 0.806, improving to 0.822 after merging. ResNet 110 models had baseline accuracies between 0.744 and 0.814, with the merged model achieving 0.824. For ResNet 152 models, baseline accuracies ranged from 0.747 to 0.783, and the merged model reached 0.815. Despite the larger complexity of ResNet 152, our method effectively enhanced model performance.